OpenStudy (anonymous):

A person pushes a shopping cart with a force of 98 N acting down at 45°. If the cart travels at a constant speed of 0.8 m/s, what is the value of the combined retarding force acting horizontally on the cart due to friction and air drag? [Hint: It is the horizontal component of the force that drives the cart forward. The constant speed tells you that Fx = ma = 0.]

OpenStudy (anonymous):

The value of the combined retarding force acting horizontally on the cart due to friction and air drag must be equal to de X component of the force 98N. \[\sum_{?}^{?}F= m*a=0\] so \[Fx- Fair-Ffriction=0\] then: Cos(45)* 98N = F air + F friction = 0,6995 N (action equals minus reaction)

OpenStudy (anonymous):

I am still confused. What's the final answer? and why?

OpenStudy (anonymous):

Nevermind I got it :] thanks